OpenStudy (anonymous):

Solve for w k=2b-5w/3k

OpenStudy (alexandervonhumboldt2):

3k^2=2b-5w 3k^2-2b=-5w -(3k^2-2b)/5=w

OpenStudy (anonymous):

Thank you, I understand how you got it